OpenStudy (anonymous):

Do the lines x-3y+4=0 and 6x-18y+24=0 have a common direction vector of (3,1)? I get (3,1) but my textbook says it's (3,2)

OpenStudy (nincompoop):

show your solution here

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 3 }x+\frac{ 4 }{ 3 }\] This is the equation that I got for both lines. And since \[m=\frac{ 1 }{ 3 }\] the direction vector m=(3,1)

OpenStudy (amistre64):

x -3y +4 =0 6x -18y +24=0 line 2 is just a multiple of line 1 1/6 = 3/18 = 4/24 so yes, they have the same slope/ same direction vector: 1/3

OpenStudy (amistre64):

your text is either misprinted, or youve got an error in the way your problem looks